# Magic Boundaries of 3D Color Codes

### v3

@@ III.3.3 The folded boundaries
[…]
III.3.4 The $pg_{\mathbf{x}}py_{\mathbf{x}}yg_{\mathbf{x}}$-boundary
The condensation set for the $pg_{\mathbf{x}}py_{\mathbf{x}}yg_{\mathbf{x}}$-boundary is given by $\{y_{\mathbf{z}}g_{\mathbf{z}},g_{\mathbf{z}}p_{\mathbf{z}},pg_{\mathbf{x}}py_{\mathbf{x}}yg_{\mathbf{x}}\}$. Similar to the folded-boundary cases, we need to couple ancilla qubits to the system to make the boundary stabilizers be mutually commuting. In general, we need to have one ancilla qubit for every pair of $\mathcal{A}^{boundary}_{y}(X)$, $\mathcal{A}^{boundary}_{g}(X)$ and $\mathcal{A}^{boundary}_{p}(X)$ stabilizers, as suggested in the 3D toric code model example in Appx. B.2.4. However, because of the lattice is asymmetry between different colors, the numbers of $X$-stabilizers for each colors don’t match. To solve this problem, we need to change the basis of generators for the $\mathcal{A}^{boundary}_{p}(X)$ and $\mathcal{B}^{boundary}_{yg}(Z)$ stabilizers, which are given by
[…]
One can check they anti-commute with each other, while commute with all the other stabilizers. Similar to the cases for the folded boundaries, we introduce ancilla qubits and couple them with the stabilizers. We introduce three ancilla qubits per set of $X$-stabilizers. Each set of $X$-stabilizers contains one $X$-stabilizer from each color. For example, we add ancilla qubits as follows:
@@ III.3.4 The $pg_{\mathbf{x}}py_{\mathbf{x}}yg_{\mathbf{x}}$-boundary
One can check those terms are mutually commuting, and commute with the string or membrane operators that generate corresponding types of excitations which can condense on it. This boundary is equivalent to the $(e_{1}e_{2},e_{2}e_{3},m_{1}m_{2}m_{3})$-boundary after we applying the unfolding unitaries. Namely,
@@
The total number of distinct types of $pg_{\mathbf{x}}py_{\mathbf{x}}yg_{\mathbf{x}}$-boundaries is 1. A 3D toric code version of this boundary is constructed in Appx. B.2.4.
III.3.5 The $y_{\mathbf{z}}g_{\mathbf{z}}p_{\mathbf{z}}$-boundary
The condensation set for the $y_{\mathbf{z}}g_{\mathbf{z}}p_{\mathbf{z}}$-boundary is given by $\{pg_{\mathbf{x}}py_{\mathbf{x}},py_{\mathbf{x}}yg_{\mathbf{x}},y_{\mathbf{z}}g_{\mathbf{z}}p_{\mathbf{z}}\}$. To construct the lattice Hamiltonian, we do the same trick as Eq. (27) to match the number of stabilizers. However, in this case we couple one ancilla qubit per $X$-stabilizer set, as suggested in the 3D toric code model example in Appx. B.2.5. The ancilla qubits are placed as follows:
@@ III.3.5 The $y_{\mathbf{z}}g_{\mathbf{z}}p_{\mathbf{z}}$-boundary
The total number of distinct types of $y_{\mathbf{z}}g_{\mathbf{z}}p_{\mathbf{z}}$-boundaries is 1. A 3D toric code version of this boundary is constructed in Appx. B.2.5.
III.3.6 Other boundaries with composite condensations
@@
III.3.7 Summary of the elementary boundary types
@@
IV Boundaries from attaching domain walls
@@
This boundary is not equivalent to the original all-smooth boundary.
We briefly review the result of Ref. [13] in this section and establish the equivalence of this boundary and the magic boundary we propose in this paper.
IV.1 Domain-wall sweeping
@@
In this case, we say $\widetilde{U}$ corresponds to a certain type of emergent symmetries for the system, including 0-form symmetries which act on the entire system and higher-form ($q$-form) symmetries which act on a codimension-$q$ submanifold [^5] Here we only consider the spatial dimensions. Note that a codimension-$q$ submanifold of a $d$-dimensional system corresponds to a dimension-$(d-q)$ submanifold. An important property of a higher-form symmetry operator is that it is equivalent up to local deformation as long as its support, the codimension-$q$ submanifold, belongs to the same homotopy class.
@@
We denote the code space of the boundary of the system as $\mathfrak{B}$. To satisfy the condition in Eq. (41), the transversal gate $\widetilde{U}$ must keep the boundary code space invariant. We have
@@
IV.2 An overview of domain walls and boundaries using the TQFT description
@@ IV.2.1 The $(e_{i},m_{j},m_{k})$-boundary
where the $\equiv$’ in the second line comes from the fusion rule is due to the fact that two invertible domain walls fuse into vacuum $s^{(2)}_{i,k}\times s^{(2)}_{i,k}=\mathbb{I}$ and hence $m_{j}s^{(2)}_{i,k}$$\times$$s^{(2)}_{i,k}=m_{j}$. Therefore, the $s^{(3)}_{1,2,3}$-domain wall also condenses on the $(e_{i},m_{j},m_{k})$-boundary.
@@
$s_{i,j}^{(2)}:(e_{i},m_{j},m_{k})\to(e_{i},m_{j}e_{i},m_{k}).$ (52)
Here, a codimension-2 (1D) boundary $(e_{i},m_{j}e_{i},m_{k})$ is produced at the interface of two codiemsnion-1 (2D) boundaries $(e_{i},m_{j},m_{k})$. We dub this type of boundaries as the nested boundaries, as illustrated in Fig. 9(a) in the later subsection.
@@ IV.2.2 The $(e_{i},e_{j},m_{k})$- and $(e_{1},e_{2},e_{3})$-boundaries
For the $(e_{1},e_{2},e_{3})$-boundaries, the condensation follows obviously from Eq. (48) and Eq. (IV.2) since both $s^{(2)}_{i,j}$ and $s^{(3)}_{1,2,3}$ act trivially on the $e$-particles.
IV.2.3 The folded boundaries
@@
$\to$ $(e_{i}e_{j},m_{i}m_{j},m_{k}e_{i}e_{j})\equiv(e_{i}e_{j},m_{i}m_{j},m_{k}),$
which shows that the composite of domain walls, $s_{j,k}^{(2)}s_{i,k}^{(2)}$, can condense on the boundary. We further have
@@
$\equiv(e_{i}e_{j},m_{i}m_{j},m_{k})$
Therefore, the $s_{i,j,k}^{(3)}$ domain wall can condense on both types of the folded boundaries.
@@ IV.2.5 The $(m_{1},m_{2},m_{3})$-boundary
IV.2.6 Other boundaries classified by the Lagrangian subgroup
@@
IV.3 Transversal gates on the 3D color code
@@
which leads to $P_{C}\widetilde{T}H^{bulk}\widetilde{T}^{\dagger}P_{C}$$=$$P_{C}H^{bulk}(X)P_{C}$ and satisfing the condition in Eq. (44). From the symmetry perspective, this is equivalent to say that transversal-$T$ gate corresponds to an emergent 0-form symmetry of the code space in the bulk.
@@
which is illustrated in Fig. 6(b) and corresponds to the mapping Eq. (IV.2) in the TQFT language.
@@
In the subsequent subsection, we demonstrate explicitly that this boundary corresponds to the magic boundary previously outlined in the context of TQFT, as illustrated in Sec.IV.2. Therefore, we also refer to this new boundary as the magic boundary. The origin of this name is due to the fact that the Hamiltonian of this boundary is beyond the Pauli stabilizers formalism, and hence possess ‘magic[37]. The magic boundary is also an important ingredient for performing fault-tolerant non-Clifford logical gates in fractal codes which can replace the usual magic state distillation protocol as pointed out in Ref. [13]. This magic boundary can also be viewed as a composite boundary, comprising the $X$-stabilizers on the boundaries attached with $S$-domain walls (equivalent to $s^{(3)}_{1,2,3}$), as depicted in Fig. 7(a).
IV.4 The magic boundary
@@
The magic boundary is exotic since the condensations are beyond the Lagrangian subgroup. One can see it by observing neither $X$-type nor $Z$-type excitations can condense on it.
@@
Here, we utilize gray circles to denote the intersecting vertices between $\widetilde{X}_{py}$ and $\mathcal{A}^{bulk}_{y}(XS)$. As shown by Eq. (78), upon exchanging the order of $\widetilde{X}_{py}$ and $\mathcal{A}^{bulk}_{y}(XS)$, an additional $Z$-operator, accompanied by a phase factor of $i$ or $-i$, emerges at each gray circle. These phase factors cancel themselves in our convention.
@@
However, in the presence of boundaries, translation symmetry is broken, and the stabilizers on the boundaries may not be identical to those in the bulk. Consequently, after conjugating transversal-$T$ gate on the entire system, $\widetilde{X}_{\mathbf{c}_{i}\mathbf{c}_{j}}$ cannot terminate on the magic boundary, namely, the $\{pg_{\mathbf{xs}},py_{\mathbf{xs}},yg_{\mathbf{xs}}\}$-boundary.
We follow the same procedure to check the commutation relations on the magic boundary. Define $\mathcal{A}^{boundary}_{\mathbf{c}_{i}}(XS)$ to be the $XS$-stabilizers on the truncated $\mathbf{c}_{i}$-cells on the boundary, which we introduce in Sec. III.2. We first consider the commutator between $\widetilde{X}_{py}$, as shown in Fig. 4(b), and $\mathcal{A}^{boundary}_{y}(XS)$, which is given by
@@
Furthermore, since the $p_{\mathbf{z}},y_{\mathbf{z}}$, and $g_{\mathbf{z}}$ excitations do not condense on the $X$-boundary, as we argued in Sec. III.2, it is straightforward to check they do not condense on the magic boundary as well.
@@
IV.5 The other elementary boundaries
@@
IV.5.1 The $\{(\mathbf{c}_{i})_{\mathbf{z}},(\mathbf{c}_{i}\mathbf{c}_{k})_{\mathbf{x}},(\mathbf{c}_{i}\mathbf{c}_{j})_{\mathbf{x}}\}$-boundaries
@@
Without loss of generality, let us consider the boundary that condenses $\{y_{\mathbf{z}},py_{\mathbf{x}},yg_{\mathbf{x}}\}$, as we introduce in Sec. III.3.1.
@@
We follow the same procedure in Sec. IV.4. The commutator between $\overline{X}_{py}$ and $\mathcal{A}^{boundary}_{p}(XS)$ is the following
@@
which is equivalent to Eq. (IV.2.1) in the TQFT description in Sec. IV.2. Although the 3D color code has color permutation symmetry, by purely permuting the color labels, the boundaries we get are physically equivalent. Therefore, we conclude that there are 3 distinct types of boundaries which can condense one $Z$-type excitations on them. For the 3D color code with these type of boundaries, the transversal-$T$ gate corresponds to an emergent 0-form symmetry of the entire code space.
@@
in which $\mathbf{c}_{i},\mathbf{c}_{j},\mathbf{c}_{k}\in\mathbf{C}$ and $i\neq j\neq k$. This general relation also corresponds to Eq. (IV.2.1).
@@
IV.5.2 The $\{(\mathbf{c}_{i})_{\mathbf{z}},(\mathbf{c}_{j})_{\mathbf{z}},(\mathbf{c}_{i}\mathbf{c}_{j})_{\mathbf{x}}\}$-boundary
Without loss of generality, we consider the $\{y_{\mathbf{z}},g_{\mathbf{z}},yg_{\mathbf{x}}\}$-boundary to illustrate the properties of this type of boundary. The boundary stabilizers can be found in Sec. III.3.2.
@@
IV.5.3 The $\{y_{\mathbf{z}},g_{\mathbf{z}},p_{\mathbf{z}}\}$-boundary

### v4

@@ 3.3.3 The folded boundaries
One can check the stabilizers above are mutually commuting, and commute with the string or membrane operators that generate corresponding types of excitations which can condense on it. The total number of distinct types of fold|e|e-boundaries is 3.
We then consider the $\{y_{\mathbf{z}}g_{\mathbf{z}},pg_{\mathbf{x}}py_{\mathbf{x}},yg_{\mathbf{x}}\}$-boundary as an example for the fold|m|m-boundary. The stabilizers on this boundary are similar to those of the previous type, except $\mathcal{B}^{boundary}_{yg}(Z)$ is substituted with $\mathcal{A}^{boundary}_{p}(X)$, which is
@@
Therefore, the generating set of stabilizers on this boundary consists of the first stabilizer in Eq. (24), the stabilizers in Eq. (25), and the rightmost stabilizer in Eq. (26). The total number of distinct types of fold|m|m-boundaries is 3.
3.3.4 The $pg_{\mathbf{x}}py_{\mathbf{x}}yg_{\mathbf{x}}$-boundary
The condensation set for the $pg_{\mathbf{x}}py_{\mathbf{x}}yg_{\mathbf{x}}$-boundary is given by $\{y_{\mathbf{z}}g_{\mathbf{z}},g_{\mathbf{z}}p_{\mathbf{z}},pg_{\mathbf{x}}py_{\mathbf{x}}yg_{\mathbf{x}}\}$. Similar to the folded-boundary cases, we need to couple ancilla qubits to the system to make the boundary stabilizers be mutually commuting. In general, we need to have one ancilla qubit for every pair of $\mathcal{A}^{boundary}_{y}(X)$, $\mathcal{A}^{boundary}_{g}(X)$ and $\mathcal{A}^{boundary}_{p}(X)$ stabilizers, as suggested in the 3D toric code model example in Appendix B.2.4. However, due to the asymmetry of the lattice with respect to different colors, the numbers of $X$-stabilizers for each colors don’t match. To solve this problem, we need to change the basis of generators for the $\mathcal{A}^{boundary}_{p}(X)$ and $\mathcal{B}^{boundary}_{yg}(Z)$ stabilizers, which are given by
@@ 3.3.4 The $pg_{\mathbf{x}}py_{\mathbf{x}}yg_{\mathbf{x}}$-boundary
One can check they anti-commute with each other, while commute with all the other stabilizers. Similar to the cases for the folded boundaries, we introduce ancilla qubits and couple them with the stabilizers. We introduce three ancilla qubits per set of $X$-stabilizers. Each set of $X$-stabilizers contains one $X$-stabilizer from each color. For the explicit forms, one may refer to Eq. (25) and the stabilizer shown on the left side of Eq. (27). For example, we add ancilla qubits as follows:
@@
One can check those terms in Eq. (29-32) are mutually commuting, and commute with the string or membrane operators that generate corresponding types of excitations which can condense on it. This boundary is equivalent to the $(e_{1}e_{2},e_{2}e_{3},m_{1}m_{2}m_{3})$-boundary after we applying the unfolding unitaries. Namely,
@@
The total number of distinct types of $pg_{\mathbf{x}}py_{\mathbf{x}}yg_{\mathbf{x}}$-boundaries is 1. A 3D toric code version of this boundary is constructed in Appendix B.2.4.
3.3.5 The $y_{\mathbf{z}}g_{\mathbf{z}}p_{\mathbf{z}}$-boundary
The condensation set for the $y_{\mathbf{z}}g_{\mathbf{z}}p_{\mathbf{z}}$-boundary is given by $\{pg_{\mathbf{x}}py_{\mathbf{x}},py_{\mathbf{x}}yg_{\mathbf{x}},y_{\mathbf{z}}g_{\mathbf{z}}p_{\mathbf{z}}\}$. To construct the lattice Hamiltonian, we do the same trick as Eq. (27) to match the number of stabilizers. However, in this case we couple one ancilla qubit per $X$-stabilizer set, as suggested in the 3D toric code model example in Appendix B.2.5. The ancilla qubits are placed as follows:
@@ 3.3.5 The $y_{\mathbf{z}}g_{\mathbf{z}}p_{\mathbf{z}}$-boundary
The total number of distinct types of $y_{\mathbf{z}}g_{\mathbf{z}}p_{\mathbf{z}}$-boundaries is 1. A 3D toric code version of this boundary is constructed in Appendix B.2.5.
3.3.6 Other boundaries with composite condensations
@@
3.3.7 Summary of the elementary boundary types
@@
4 Boundaries from attaching domain walls
@@
$s_{i,j}^{(2)}$ is a line defect, whose definition and detailed discussion can be found in Ref. [13]. This boundary is not equivalent to the original all-smooth boundary.
The structure of this section is organized as follows. In subsection 4.1, we review the domain wall sweeping picture for implementing (generalized) global symmetries. In subsection 4.2, we discuss the classification of boundaries in $\mathbb{Z}_{2}^{3}$ gauge theories using the language of TQFT, which follows Ref. [13]. In subsection 4.3, we construct the explicit form of transversal gates on the 3D color code defined on our chosen lattice, and analyze their transformation under the application of transversal $T$ and $S$ gates. In subsection 4.4, we study the action of logical operators on the $T$-conjugated ground state. In subsection 4.5, we construct the boundary Hamiltonian for the magic boundary and study its exotic properties. In subsection 4.6, we construct the boundary Hamiltonians for other types of elementary boundaries. Finally, in subsection 4.7, we study the nested boundaries, which are boundaries that have codimension 1 condensation defects.
4.1 Domain wall sweeping
@@
In this case, we say $\widetilde{U}$ corresponds to a certain type of emergent symmetries for the system, including 0-form symmetries which act on the entire system and higher-form ($q$-form) symmetries which act on a codimension-$q$ submanifold [^5] Here we only consider the spatial dimensions. Note that a codimension-$q$ submanifold of a $d$-dimensional system corresponds to a dimension-(dq)(d-q) submanifold. An important property of a higher-form symmetry operator is that it is equivalent up to local deformation as long as its support, the codimension-$q$ submanifold, belongs to the same homotopy class.
@@
We denote the code space of the boundary of the system as $\mathfrak{B}$, defined as the code space stabilized by the Hamiltonian terms supported entirely on the boundary. To satisfy the condition in Eq. (41), the transversal gate $\widetilde{U}$ must keep the boundary code space invariant. We have
@@
4.2 An overview of domain walls and boundaries using the TQFT description
@@ 4.2.1 The $(e_{i},m_{j},m_{k})$-boundary
[…]
@@
$s_{j,k}^{(2)}:(e_{i},m_{j},m_{k})\to(e_{i},m_{j}e_{k},m_{k}e_{j}).$ (52)
Here, a codimension-2 (1D) boundary $(e_{i},m_{j}e_{k},m_{k}e_{j})$ is produced at the interface of two codiemsnion-1 (2D) boundaries $(e_{i},m_{j},m_{k})$. We dub this type of boundaries as the nested boundaries, as illustrated in Fig. 9(a) in the later subsection.
@@ 4.2.2 The $(e_{i},e_{j},m_{k})$- and $(e_{1},e_{2},e_{3})$-boundaries
For the $(e_{1},e_{2},e_{3})$-boundaries, the condensation follows obviously from Eq. (48) and Eq. (4.2) since both $s^{(2)}_{i,j}$ and $s^{(3)}_{1,2,3}$ act trivially on the $e$-particles.
4.2.3 The folded boundaries
@@
$\to$ $(e_{i}e_{j},m_{i}m_{j},m_{k}e_{i}e_{j})\equiv(e_{i}e_{j},m_{i}m_{j},m_{k}).$
Here, we first apply the $s_{i,k}^{(2)}$ defect on the boundary, followed by the $s_{j,k}^{(2)}$ defect. The result shows that the composite of domain walls, $s_{j,k}^{(2)}s_{i,k}^{(2)}$, can condense on the boundary. Therefore, we have
@@
$\equiv(e_{i}e_{j},m_{i}m_{j},m_{k}).$
The $s_{i,j,k}^{(3)}$ domain wall can condense on both types of the folded boundaries.
@@ 4.2.5 The $(m_{1},m_{2},m_{3})$-boundary
4.2.6 Other boundaries classified by the Lagrangian subgroup
@@
4.3 Transversal gates on the 3D color code
@@
which leads to $P_{C}\widetilde{T}H^{bulk}\widetilde{T}^{\dagger}P_{C}$$P_{C}H^{bulk}(X)P_{C}$ and satisfing the condition in Eq. (44). From the symmetry perspective, this is equivalent to say that transversal-$T$ gate corresponds to an emergent 0-form symmetry of the code space in the bulk.
@@
which is illustrated in Fig. 6(b) and corresponds to the mapping Eq. (4.2) in the TQFT language.
@@
[…]
4.4 Logical operators after applying the transversal $T$-gate.
@@
Here, we utilize gray circles to denote the intersecting vertices between $\widetilde{X}_{py}$ and $\mathcal{A}^{bulk}_{y}(XS)$. As shown by Eq. (78), upon exchanging the order of $\widetilde{X}_{py}$ and $\mathcal{A}^{bulk}_{y}(XS)$, an additional $Z$-operator, accompanied by a phase factor of $i$ or i-i, emerges at each gray circle. These phase factors cancel themselves in our convention.
@@
4.5 The magic boundary
In the presence of boundaries, the translation symmetry is broken, and the stabilizers on the boundaries may not be identical to those in the bulk. Consequently, after conjugating transversal-$T$ gate on the entire system, $\widetilde{X}_{\mathbf{c}_{i}\mathbf{c}_{j}}$ cannot terminate on the magic boundary, namely, the $\{pg_{\mathbf{xs}},py_{\mathbf{xs}},yg_{\mathbf{xs}}\}$-boundary.
We follow the same procedure to check the commutation relations on the magic boundary. Define $\mathcal{A}^{boundary}_{\mathbf{c}_{i}}(XS)$ to be the $XS$-stabilizers on the truncated $\mathbf{c}_{i}$-cells on the boundary, which we introduce in Section 3.2. We first consider the commutator between $\widetilde{X}_{py}$, as shown in Fig. 4(b), and $\mathcal{A}^{boundary}_{y}(XS)$, which is given by
@@ 4.5 The magic boundary
Furthermore, since the $p_{\mathbf{z}},y_{\mathbf{z}}$, and $g_{\mathbf{z}}$ excitations do not condense on the $X$-boundary, as we argued in Section 3.2, it is straightforward to check they do not condense on the magic boundary as well.
@@
4.6 The other elementary boundaries
@@
4.6.1 The $\{(\mathbf{c}_{i})_{\mathbf{z}},(\mathbf{c}_{i}\mathbf{c}_{k})_{\mathbf{x}},(\mathbf{c}_{i}\mathbf{c}_{j})_{\mathbf{x}}\}$-boundaries
@@
Without loss of generality, let us consider the boundary that condenses $\{y_{\mathbf{z}},py_{\mathbf{x}},yg_{\mathbf{x}}\}$, as we introduce in Section 3.3.1.
@@
We follow the same procedure in Section 4.5. The commutator between $\overline{X}_{py}$ and $\mathcal{A}^{boundary}_{p}(XS)$ is the following
@@
which is equivalent to Eq. (4.2.1) in the TQFT description in Section 4.2. Although the 3D color code has color permutation symmetry, by purely permuting the color labels, the boundaries we get exhibit identical physical properties, differing only by relabeling of the lattice and excitations. Therefore, we conclude that there are 3 distinct types of boundaries which can condense one $Z$-type excitations on them. For the 3D color code with these type of boundaries, the transversal-$T$ gate corresponds to an emergent 0-form symmetry of the entire code space.
@@
in which $\mathbf{c}_{i},\mathbf{c}_{j},\mathbf{c}_{k}\in\mathbf{C}$ and $i\neq j\neq k$. This general relation also corresponds to Eq. (4.2.1).
@@
4.6.2 The $\{(\mathbf{c}_{i})_{\mathbf{z}},(\mathbf{c}_{j})_{\mathbf{z}},(\mathbf{c}_{i}\mathbf{c}_{j})_{\mathbf{x}}\}$-boundary
Without loss of generality, we consider the $\{y_{\mathbf{z}},g_{\mathbf{z}},yg_{\mathbf{x}}\}$-boundary to illustrate the properties of this type of boundary. The boundary stabilizers can be found in Section 3.3.2.
@@
4.6.3 The $\{y_{\mathbf{z}},g_{\mathbf{z}},p_{\mathbf{z}}\}$-boundary